Victory Ltd commences operations on 1 July 2020. One year later, on 30 June 2021, the entity prepares its first statement of comprehensive income and its first statement of financial position. The statements are prepared before considering taxation. The following information is available. Statement of comprehensive income for the year ended 30 June 2021 Gross profit Wages expense Long service leave expense Bad debts expense Rent expense Depreciation expense machinery Accounting profit before tax S 500000 (200000) 50000) (20000) (50000) (30000) 150000 Assets and liabilities as disclosed in the statement of financial position as at 30 June 2021 Assets $150000 200000 180000 50000 150000) (30000 700000 Accounts receivables (net) Prepaid rent Machinery Accounts payable Revenue received in advance Loan payable Provision for long service leave 100000 50000 200000 50000 400000 Additional information The company tax rate is assumed to be 30%. . All salaries have been paid as at year end and are deductible for tax purposes. . None of the long service leave expense has actually been paid. It is not deductible for . Rent was paid in advance on 1 July 2020. Actual amounts paid are allowed as a tax . Amounts received from sales, including those on credit terms, are taxed at the time . The revenue received in advance is included in the taxable income. tax purposes until it is actually paid. deduction. the sale is made. No bad debts were written off. . The machinery is depreciated on a straight-line basis over 5 years for accounting purposes, but over 3 years for taxation purposes. The machinery is not expected to have any residual value.
How to we determine the future deductible amount for machinery? ($100,000)
Answer #1

Depreciation on Straight line method

Straight line depreciation is the default method used to gradually reduce the carrying amount of a fixed asset over its useful life. The Straight line method is the easiest depreciation method to calculate.

Formula

Depreciation per annum = Cost - Residual value or salvage value or spare value

Useful life

or

Depreciation per annum = (Cost - Residual value or salvage value or spare value) x Rate of Depreciation

So as per the above question the calculation as follows:

Depreciation for Accounting Purpose:

Cost of Machinery = 150000

Residual Value = Nil

Life = 5 year

Forumula:

Dep per annum = Cost - Residual Value / Useful life

= 150000-0 / 5

Depreciation Exp = 30000

Hence Depreciation will be charged as per the accounting purpose for next remaining 4 years 30000 p.a.

1st year Dep = 150000-30000 = 120000

2nd year Dep = 120000-30000 = 90000

3rd year Dep = 90000-30000 = 60000

4th year Dep = 60000-30000 = 30000

5th year Dep = 30000-30000 = Nil

Depreciation for Taxation Purpose:

Cost of Machinery = 150000

Residual Value = Nil

Life = 3 year

Forumula:

Dep per annum =   Cost - Residual Value / Useful life

=   150000-0 / 3

Depreciation Exp = 50000

Hence Depreciation will be charged as per the accounting purpose for next remaining 2 years 50000 p.a.

1st year Dep = 150000-50000 = 100000

2nd year Dep = 100000-50000 = 50000

3rd year Dep = 50000-50000 = Nil

So the future deductible amount for machinery as per the Accounting purpose would be 30000

and the Taxation purpose would be 50000